# Equivalence principles in Weyl transverse gravity

### v1

@@ Appendix A The WTDiff-invariant Ehlers criterion
In this appendix, we provide a WTDiff-invariant formulation of one of the well-established criteria for matter fields compatible with the Einstein EP, the Ehlers criterion [40] (using its “g-approximating” version proposed in [4]). For a WTDiff-invariant theory, the criterion can be stated in the following way. We start with a manifold $M$ endowed with a dynamical metric $g_{\mu\nu}$ and a nondynamical volume $D$-form $\bm{\omega}$ (which together allow us to construct the auxiliary metric $\tilde{g}_{\mu\nu}$). We further consider a collection of fields $\psi$. We prescribe the equations of motion for the matter fields to be $A[\tilde{g}_{\mu\nu},\psi]=0$, where $A[\tilde{g}_{\mu\nu},\psi]$ is a differential expression constructed from the auxiliary metric $\tilde{g}_{\mu\nu}$, the fields $\psi$, and their derivatives. In a neighbourhood $O$ of an arbitrary regular point $P\in M$ we introduce a flat spacetime metric $\eta_{\mu\nu}$ such that it agrees with $\tilde{g}_{\mu\nu}$ in any point $Q\in O$ to the first order in the geodesic distance between $P$ and $Q$. Consider any $\bar{\psi}$ that solves the equation $A[\eta_{\mu\nu},\bar{\psi}]=0$ in $O$. Then, for any $\epsilon>0$, there exists some $\psi$ solving the curved spacetime equation $A[\tilde{g}_{\mu\nu},\psi]=0$ in some neighborhood $O^{\prime}$ of $P$, such that $O^{\prime}\subseteq O$, which approximates $\bar{\psi}$. More precisely, we require that there exists some neighborhood $U$ of $P$, such that $U\subseteq O^{\prime}$ and $||\bar{\psi}-\psi||<\epsilon$ in $U$. The choice of a suitable norm |||||$\cdot$||| | ⋅ | | depends on the context, one possible example being the $C^{k}$ supremum norms on the sub-neighborhood $U$ [4].

### v2

@@ Appendix A The WTDiff-invariant Ehlers criterion
In this appendix, we provide a WTDiff-invariant formulation of one of the well-established criteria for matter fields compatible with the Einstein EP, the Ehlers criterion [41] (using its “g-approximating” version proposed in [4]). For a WTDiff-invariant theory, the criterion can be stated in the following way. We start with a manifold $M$ endowed with a dynamical metric $g_{\mu\nu}$ and a nondynamical volume $D$-form $\bm{\omega}$ (which together allow us to construct the auxiliary metric $\tilde{g}_{\mu\nu}$). We further consider a collection of fields $\psi$. We prescribe the equations of motion for the matter fields to be $A[\tilde{g}_{\mu\nu},\psi]=0$, where $A[\tilde{g}_{\mu\nu},\psi]$ is a differential expression constructed from the auxiliary metric $\tilde{g}_{\mu\nu}$, the fields $\psi$, and their derivatives. In a neighbourhood $O$ of an arbitrary regular point $P\in M$ we introduce a flat spacetime metric $\eta_{\mu\nu}$ such that it agrees with $\tilde{g}_{\mu\nu}$ in any point $Q\in O$ to the first order in the geodesic distance between $P$ and $Q$. Consider any $\bar{\psi}$ that solves the equation $A[\eta_{\mu\nu},\bar{\psi}]=0$ in $O$. Then, for any $\epsilon>0$, there exists some $\psi$ solving the curved spacetime equation $A[\tilde{g}_{\mu\nu},\psi]=0$ in some neighborhood $O^{\prime}$ of $P$, such that $O^{\prime}\subseteq O$, which approximates $\bar{\psi}$. More precisely, we require that there exists some neighborhood $U$ of $P$, such that $U\subseteq O^{\prime}$ and $\bar{\psi}-\psi||<\epsilon$ in $U$. The choice of a suitable norm $||\cdot||$ depends on the context, one possible example being the $C^{k}$ supremum norms on the sub-neighborhood $U$ [4].